OpenStudy (anonymous):

What is the value of the expression? (–4)4

OpenStudy (anonymous):

Soo 4*-4?

OpenStudy (anonymous):

-4*4=-16

OpenStudy (mrnood):

is this meant to be \[-4^{4}\]

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

For future reference, just use this key to show a square: ^. So it would be shown as (-4)^4

OpenStudy (anonymous):

(-4)^4

OpenStudy (mrnood):

well raise to an even power means that the negative goes away (e.g. -2 * -2 = -2^2 = 4 so this is the same as 4^4

OpenStudy (anonymous):

it = -16 right

OpenStudy (anonymous):

?

OpenStudy (anonymous):

(-4)^4=-256

OpenStudy (mrnood):

no it is not negative and it is not 16

OpenStudy (mrnood):

@cb77613 that is not correct

OpenStudy (anonymous):

sorry it is 256

OpenStudy (anonymous):

then what is it?

OpenStudy (anonymous):

forgot to put in () it is 256

OpenStudy (mrnood):

@DrumRBoy1927 look at it as -4 * -4 *-4 *-4 you know that multiplyint TWO negatives gives a positive so does multiplying FOUR negativex

OpenStudy (mrnood):

@cb77613 It is against the code of conduct to post direct answers - and it does not help the original poster either

OpenStudy (anonymous):

You guys are confusing me

OpenStudy (mrnood):

@DrumRBoy1927 have another look at it as (-4*-4) * (-4*-4) you can see it ends up as a positive

OpenStudy (anonymous):

256

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

anything else you need help with